Problem 15-5 Calculating Option Payoffs (LO2, CFA2)

Strike Calls Puts
Close Price Expiration Vol. Last Vol. Last
Hendreeks
103 100 Feb 72 5.20 50 2.40
103 100 Mar 41 8.40 29 4.90
103 100 Apr 16 10.68 10 6.60
103 100 Jul 8 14.30 2 10.10

Suppose you buy 35 April 100 call option contracts. Hendreeks stock is selling for $105.90 per share on the expiration date. How much is your options investment worth? What if the stock price is $101.80 on the expiration date? (Do not round intermediate calculations.)

Terminal value at 105.90=

Terminal value at 101.80=

----------------------------------------------------------------------------------------------------------------------------------------------------------

Strike Calls Puts
Close Price Expiration Vol. Last Vol. Last
Hendreeks
103 100 Feb 72 5.20 50 2.40
103 100 Mar 41 8.40 29 4.90
103 100 Apr 16 10.68 10 6.60
103 100 Jul 8 14.30 2 10.10

Suppose you buy 45 April 100 put option contracts. What is your maximum gain? On the expiration date, Hendreeks is selling for $85.80 per share. How much is your options investment worth? What is your net gain? (Do not round intermediate calculations.)

Maximum gain=

Terminal value=

Net gain=
Answer #1

Answers

1) Cost of 35 April 100 Call option contracts = 35 x 100 x 10.68 = $37,380

Hence the option investment is worth $37,380.

Terminal value @105.90 = 35x100x(105.90-100)= $20650

Terminal value @101.80 = 35x100x(101.80-100)= $6300

2) Cost of 45 April 100 Put option contracts = 45 x 100 x 10.68 =$48,060

Hence the option investment is worth $48,060.

Maximum gain would be possible if entire $100 of strike price is net gain per share, hence maximum gain would be

= 45x100x100 - Initial cost = $450,000-$48,060 = $401,940

Terminal value @85.80 = 45x100x(100-85.80) = $63,900

Net gain = Terminal value - Initial cost = $63900-$48,060 = $15,840